OpenStudy (anonymous):

3. Evaluate: Log4 1/64

OpenStudy (anonymous):

Do you mean: \[\log_{4} (\frac{ 1 }{ 64 })\]

OpenStudy (anonymous):

Yes exactly

OpenStudy (anonymous):

What I do to solve problems like this is set them equal to x and then use the logarithm rules and exponential rules to help solve. So: \[\log_{4} (\frac{ 1 }{ 64 })=x\] Now we can say that: \[\frac{ 1 }{ 64 }=4 ^{x}\] Now just change 1/64 so it has a base of 4 and you have your answer. Let me know if you want me to post the answer.

OpenStudy (anonymous):

log4 64=1 ? @Directrix can you help?

Directrix (directrix):

Picking up on the previous poster's work: Now just change 1/64 so it has a base of 4 and you have your answer. Let me know if you want me to post the answer.